### v1

# August 6, 2024

Abstract
The real part of pseudo entropy in conformal field theories is holographically calculated by the area of some extremal spacelike surfaces in the dual dS and AdS spacetimes. We show that the flat-space limit of these curves in three dimensional (A)dS spacetimes is well-defined. We find that if the length of the curves is calculated from the radial coordinate where the retarded time is extremum, then after taking the flat-space limit, the entanglement entropy of the dual theory of three-dimensional flat spacetime is obtained. For dS spacetime, the radial coordinate corresponding to the extremum of retarded time is located inside the cosmlogical horizon. Our results suggest that on the field theory side, the entanglement entropy in the dual theory of flat spacetimes should be obtained from the ultra-relativistic limit of pseudo entropy in the dual CFT to (A)ds spacetimes.
1 Introduction
It has been proposed in [1]-[2] that the flat-space limit (i.e., the zero cosmological constant limit) of asymptotically AdS spacetimes is equivalent to the ultra-relativistic limit [3] in conformal field theory. As a result, it has been suggested that the dual field theory to asymptotically flat spacetimes in three dimensions is a two-dimensional ultra-relativistic field theory, known as Carrollian Conformal Field Theory (CCFT) [2, 4, 5]. This theory exists at null infinity of asymptotically flat spacetimes, and this duality is referred to as the flat/CCFT correspondence.  [^1] In some papers, this duality is also called flat/BMSFT. BMSFT refers to a field theory with BMS symmetry, which is the asymptotic symmetry of asymptotically flat spacetimes at null infinity [6]-[9]. This symmetry is equivalent to Carrollian conformal symmetry in one lower dimension.
One notable work in the field of AdS/CFT is the Ryu-Takayanagi proposal. According to this proposal, the entanglement entropy in conformal field theories (CFTs) corresponds to the area of extremal surfaces in asymptotically AdS spacetimes. These extremal surfaces are connected to the subsystem on the AdS boundary for which the entanglement entropy is calculated. This correspondence is significant because it simplifies the calculation of entanglement entropy in field theory by relating it to more straightforward calculations in gravity.
Typically, a general system is decomposed into subsystems at constant time levels. This allows the Hilbert space to be expressed as a tensor product of its subspaces. From a general density matrix in the original system, one can derive a reduced density matrix for spacelike subsystems and calculate the entanglement entropy. For CFTs, the holographic dual of this entanglement entropy, according to Ryu-Takayanagi [10], is represented by extremal spacelike surfaces in the bulk of asymptotically AdS spacetimes. Recently, [11]-[12] have studied the generalization of entanglement entropy for timelike subsystems and proposed holographic duals for these as well. This entropy, known as timelike entanglement entropy, is a complex quantity where the real part corresponds to the area of spacelike extremal surfaces and the imaginary part corresponds to the area of timelike extremal surfaces in the bulk. In asymptotically AdS spacetimes, these spacelike surfaces are divided into several parts connected to the subsystem at the boundary but not directly to each other in the bulk, with timelike surfaces connecting them. Furthermore, [11]-[12] argue that timelike entanglement entropy is essentially a special case of pseudo-entropy [13], where instead of the reduced density matrix, the transition matrix is used (see section 2.1).
The Ryu-Takayanagi proposal for holographic entanglement entropy can be explored in other dualities. For asymptotically de Sitter (dS) spacetimes, this correspondence is known as dS/CFT, suggesting that a (d-1)-dimensional Euclidean conformal field theory is the holographic dual of d-dimensional asymptotically dS spacetimes [14]. This field theory, defined at the timelike infinity of dS spacetime, is non-unitary, so the entanglement entropy for its subsystems is essentially pseudo-entropy and must be complex. The holographic dual of this pseudo-entropy is given in [11]-[12], consisting of extremal surfaces in the bulk, with parts being spacelike and others timelike. Unlike in AdS, the timelike surfaces here connect to the subsystem at timelike infinity, while the spacelike surfaces appear inside the bulk and connect to the timelike surfaces.
For CCFTs, as the dual theory of asymptotically flat spacetimes, entanglement entropy can also be defined. Using the infinite-dimensional symmetry of two-dimensional CCFTs, a universal formula for the entanglement entropy of intervals is provided in [15]. The holographic interpretation of this entropy is proposed in [16] as extremal curves in asymptotically flat spacetimes. These curves consist of spacelike and null parts, with the spacelike part entirely in the bulk and the null part connecting the ends of the interval at null infinity to the spacelike curve.
The method used in [16] to find extremal surfaces generalizes Rindler’s method [17] to flat/CCFT. By applying a unitary transformation known as the Rindler transformation, entanglement entropy is mapped to thermal entropy, whose holographic counterpart is the entropy of the cosmological horizon in three-dimensional asymptotically flat spacetimes, known as flat-space cosmology (FSC) [18]-[19] (see also [20]). This spacetime is obtained from the flat-space limit of the BTZ black hole and has only one cosmological horizon. The Rindler transformation is extended to the bulk, and its inverse relates the FSC horizon to the extremal surface, whose area is proportional to the entanglement entropy of subsystems in CCFT.
Given that asymptotically flat spacetimes are derived from the flat-space limit of asymptotically (A)dS spacetimes, it should be possible to connect extremal surfaces in these spacetimes with those in asymptotically flat spacetimes. It is known that the flat-space limit of the extremal surfaces introduced by Ryu and Takayanagi for AdS spacetime is not well-defined. This issue is addressed in [21], which introduces new curves in three-dimensional AdS spacetime using the Rindler method. These curves map to the inner horizon of the BTZ black hole rather than the outer horizon. The flat limit of these curves leads to surfaces proposed in [16] as the holographic dual of entanglement entropy in CCFT.
In this paper, we demonstrate that the new curves in [21] are indeed the same as the spacelike part of the timelike entanglement entropy in the dual CFT to AdS. In other words, to obtain extremal curves in asymptotically flat spacetimes, one must use extremal curves corresponding to pseudo-entropy in CFT. We show that this interpretation holds for both AdS and dS spacetimes, and the spacelike part of the extremal curves, whose length corresponds to the real part of pseudo-entropy in the dual CFT, has a well-defined flat-space limit. Although the total length of the spacelike curves after the flat-space limit differs from the entanglement entropy of CCFT, we find that if we consider only the portion of these curves starting from the extremum of retarded time and extending to the end in the bulk, the length divided by 4G, where G is Newton’s constant, exactly gives the entanglement entropy of CCFT.
In the case of dS spacetime, the spacelike curve before the flat-space limit begins from a point outside the cosmological horizon and stretches to the minimum radius inside the horizon after crossing it. Since the extremum of retarded time is located inside the horizon, a segment of the spacelike curve that corresponds to the entanglement entropy of CCFT lies entirely within the cosmological horizon.
Our calculations in this paper are entirely within the gravitational sector of duality. The fact that the flat-space limit of curves corresponding to pseudo-entropy in dS and AdS results in curves corresponding to entanglement entropy in flat spacetimes suggests a similar connection on the field theory side. In other words, to find entanglement entropy in CCFT by taking the ultra-relativistic limit from a quantity in CFT, this quantity should be pseudo-entropy rather than entanglement entropy.
2 Preliminaries
2.1 Entanglement entropy and pseudo entropy in field theory
In a quantum field theory, suppose the Hilbert space ${\mathcal{H}}$ is factorized as ${\mathcal{H}}={\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$. This means that the total system is divided into two subsystems A and B, where B is the complement of A. These subsystems are assumed to be spacelike regions of the main system. If $\rho$ is the density matrix of the total system, then the reduced density matrix for subsystem A, denoted by $\rho_{A}$, is obtained by tracing out B as
$\rho_{A}=\text{Tr}_{B}\rho.$ (2.1)
The entanglement entropy for subsystem A, $S_{A}$, is the von Neumann entropy calculated using the reduced density matrix $\rho_{A}$, i.e.
$S_{A}=-\text{Tr}\left[\rho_{A}\log\rho_{A}\right].$ (2.2)
If |$\ket{\psi}$ and |$\ket{\varphi}$ are two pure states of the total system, we can define a reduced transition matrix as follows:
$\tau_{A}=\text{Tr}_{B}\left[\dfrac{\ket{\psi}\!\bra{\varphi}}{\bra{\varphi}\ket{\psi}}\right].$ (2.3)
Using this transition matrix, pseudo entropy $S_{A}^{(p)}$ is defined as [13]
$S_{A}^{(p)}=-\text{Tr}\left[\tau_{A}\log\tau_{A}\right].$ (2.4)
If subsystem A is a timelike region of the total system, the reduced density matrix can still be defined for it using formula (2.1), and the entanglement entropy can be obtained using formula (2.2). However, in this case, the entropy is complex-valued. This entropy is reffered to as timelike entanglement entropy. The complexity arises because the density matrix in this case is not Hermitian. In the paper [12], it is argued that the timelike entanglement entropy can be reformulated using a transition matrix. Therefore, the timelike entanglement entropy itself is considered a pseudo entropy.
If the field theory is non-unitary, then generally $\bra{\psi}\neq\ket{\psi}^{\dagger}$. Consequently, the operator $\rho=\ket{\psi}\!\bra{\psi}$ | behaves more like a transition matrix than a density matrix. Thus, the von Neumann entropy defined using this $\rho$ and its associated reduced matrices, will indeed be pseudo entropy.
2.2 Entanglement entropy in the dual theory of flat spacetimes
In paper [1]-[2] , it is proposed that the dual field theory to asymptotically flat spacetime in three dimensions is a two-dimensional CCFT, which possesses the following infinite-dimensional symmetry:
$[L_{n},L_{m}]=(n-m)L_{n+m}+\dfrac{c_{L}}{12}(n^{3}-n)\delta_{n+m,0},$
$[L_{n},M_{m}]=(n-m)M_{n+m}+\dfrac{c_{M}}{12}(n^{3}-n)\delta_{n+m,0},$
$[M_{n},M_{m}]=0.$ (2.5)
$c_{L}$ and $c_{M}$ are central charges.
This algebra is derived from the Inonu-Wigner contraction of the conformal algebra in two dimensions. This contraction represents the ultra-relativistic limit [3] of CFT, which is why the term Carrollian is used to describe this field theory [4, 5]. The main proposal of [2] is that the flat-space limit on the gravity side corresponds to the ultra-relativistic limit in the dual CFT.
It is evident from algebra (2.2) that the symmetries of two-dimensional CCFTs are infinite dimensional, similar to those of two-dimensional CFTs . This infinite-dimensional symmetry allow for the identification of universal properties within these theories. Among these features are universal formulas for entanglement entropy. Research into this aspect of two-dimensional CCFTs began in [15] , where it it was shown that, in a CCFT on the plane, if we consider the subsystem to be the interval $\left(-\dfrac{l_{u}}{2}<u<\dfrac{l_{u}}{2},\,-\dfrac{l_{\phi}}{2}<\phi<\dfrac{l_{\phi}}{2}\right)$ , the entanglement entropy is given by
$S_{EE}=\dfrac{c_{L}}{6}\log\dfrac{l_{\phi}}{\epsilon}+\dfrac{c_{M}}{6}\dfrac{l_{u}}{l_{\phi}},$ (2.6)
where $\epsilon$ is a cutoff used to regulate the interval. For CCFTs on the cylinder this formula is given by
$S_{EE}=\dfrac{c_{L}}{6}\log\left(\dfrac{2}{\epsilon}\sin\dfrac{l_{\phi}}{2}\right)+\dfrac{c_{M}}{12}l_{u}\cot\dfrac{l_{\phi}}{2}.$ (2.7)
The holographic interpretation of entanglement entropy was proposed in [16]. The algebra (2.2) on the gravity side corresponds to the asymptotic symmetry of asymptotically flat spacetimes at null infinity [8]. Consequently, dual CCFTs must be defined at null infinity, and each subsystem in these theories will correspond to intervals at null infinity as well. This implies that the extremal spacelike curves , whose length leads to the entropies (2.6) and (2.7) , do not connect the two ends of the interval directly. In fact, the proposal of [16] is that the holographic interpretation in this context involves spacelike curve whose endpoints are connected to the two ends of the interval at null infinity by two null rays.
3 Holographic pseudo entropy in AdS and its flat limit
In this section we aim to find the holographic timelike entanglement entropy for an interval on the boundary of AdS3 and then take the flat-space limit. As an example of pseudo entropy in CFT, this entropy is inherently complex-valued . The holographic dual of this entropy introduced in [11]-[12] consists of two geodesics: one spacelike and one timelike. To obtain the flay-space limit, we only need the spacelike part. We begin by considering a spacelike interval on the boundary of AdS3, compute its holographic entanglement entropy, and then perform a Wick rotation to extract the spacelike part of the holographic timelike entanglement entropy.
3.1 Holographic pseudo entropy from holographic entanglement entropy
Let us consider $AdS_{3}$ written in the Poincare coordinate,
$ds^{2}=\dfrac{\ell^{2}}{z^{2}}\left(-dt^{2}+dx^{2}+dz^{2}\right),$ (3.1)
where $\ell$ is the AdS radius and the boundary is located at $z=0$. First, we compute the holographic entanglement entropy for a spacelike interval in the dual CFT2 residing on the boundary of this geometry. This interval is characterized by
$-\dfrac{l_{x}}{2}<x<\dfrac{l_{x}}{2},\qquad-\dfrac{l_{t}}{2}<t<\dfrac{l_{t}}{2}$ (3.2)
where $l_{x}$ and $l_{t}$ are two constants with $l_{x}>l_{t}$, ensuring that this interval is spacelike. The entanglement entropy of this interval as a subsystem of the boundary CFT, is given by a universal formula that depends only on the length of it and the central charges of the CFT [22]. According to the Ryu-Takayanagi proposal [10], this entanglement entropy has a holographic description as the length of an extremal curve in the bulk spacetime, anchored to the two ends of the interval. In our case, this extremal curve is a semicircle given by
$x^{2}=\dfrac{l_{x}^{2}\left(l_{x}^{2}-l_{t}^{2}-4z^{2}\right)}{4\left(l_{x}^{2}-l_{t}^{2}\right)},$
$t^{2}=\dfrac{l_{t}^{2}\left(l_{x}^{2}-l_{t}^{2}-4z^{2}\right)}{4\left(l_{x}^{2}-l_{t}^{2}\right)}.$ (3.3)
Our main goal is to take the flat-space limit, $\ell\to\infty$, from the calculations and compare the results with those known for asymptotically flat spacetimes. However, the Poincare coordinate (3.1) is not suitable for this limit. Therefore, we write the AdS3 metric in global BMS coordinate [23] as
$ds^{2}=-\left(1+\dfrac{r^{2}}{\ell^{2}}\right)du^{2}-2dudr+r^{2}d\phi^{2},$ (3.4)
which is given by the following transformation from the Poincare coordinate (3.1):
$z$ $=\dfrac{2\ell^{2}}{\Delta},$
[…]
$x$ $=\dfrac{2\ell r\sin{\phi}}{\Delta},$
[…]
In this coordinate system, $u$ is known as the retarded time . If we denote the boundary interval by $-\dfrac{l_{u}}{2}<u<\dfrac{l_{u}}{2}$ and $-\dfrac{l_{\phi}}{2}<\phi<\dfrac{l_{\phi}}{2}$ then we find that
[…]
Now we can use (3.1) and (3.1) to write (3.1) in the global BMS coordinate (3.4). After some simplification, we obtain the following equations:
$r=\dfrac{\ell\cos\left(\dfrac{l_{\phi}}{2}\right)\sin\left(\frac{u}{\ell}\right)}{\cos\left(\dfrac{l_{u}}{2\ell}\right)\cos(\phi)-\cos\left(\dfrac{l_{\phi}}{2}\right)\cos(\dfrac{u}{\ell})},$
$r^{2}\sin^{2}(\phi)\sin^{2}\left(\dfrac{l_{u}}{2\ell}\right)=\sin^{2}\left(\dfrac{l_{\phi}}{2}\right)\left(\ell\cos\left(u\over\ell\right)-r\sin\left(\dfrac{u}{\ell}\right)\right)^{2}.$ (3.7)
It is clear that the flat-space limit ( $\ell\to\infty$ ) is not well-defined for the second equation (3.1). Therefore, the flat-space limit from the extremal curves given by using the Ryu-Takayanagi method for the spacelike interval is not well-defined. This implies that the holographic description of the entanglement entropy of CCFT2 (introduced in [16]) can not be obtained by taking the flat-space limit of (3.1). Our main goal is to find appropriate extremal curves with a well-defined flat-space limit. The proposal is that these curves are the holographic dual of timelike entanglement entropy. Precisely , the extremal curve with a well-defined flat space limit is obtained by interchanging ł$\ell\l_{\phi}\leftrightarrow\l_{u}$ in (3.1). The equations describing our new extremal curve are
$r=\dfrac{\ell\cos\left(\dfrac{l_{u}}{2\ell}\right)\sin\left(\frac{u}{\ell}\right)}{\cos\left(\dfrac{l_{\phi}}{2}\right)\cos(\phi)-\cos\left(\dfrac{l_{u}}{2\ell}\right)\cos(\dfrac{u}{\ell})},$
$r^{2}\sin^{2}(\phi)\sin^{2}\left(\dfrac{l_{\phi}}{2}\right)=\sin^{2}\left(\dfrac{l_{u}}{2\ell}\right)\left(\ell\cos\left(u\over\ell\right)-r\sin\left(\dfrac{u}{\ell}\right)\right)^{2}.$ (3.8)
This new curve is also spacelike and intersects the boundary at two points $\left(u=\pm\frac{\ell l_{\phi}}{2},\phi=\pm\frac{l_{u}}{2\ell}\right)$. These points mark the ends of a timelike interval on the boundary of AdS3. This interval represents the timelike (Wick-rotated) counterpart of the original spacelike interval considered as the subsystem in the dual CFT. Comparing (3.1) with the results of [11]-[12] indicates that our new curve is precisely the spacelike part of the holographic timelike entropy. The flat-space limit of this curve is
$r=\dfrac{u}{\cos\phi\cos\left({l_{\phi}\over 2}\right)-1},$
$r^{2}\sin^{2}(\phi)\sin^{2}\left(\frac{l_{\phi}}{2}\right)=\dfrac{l_{u}^{2}}{4}.$ (3.9)
which is consistent with the results of [16].
The equations of new curve (3.1), can be rewritten as
$\cos\phi=\dfrac{\cos\left(\dfrac{l_{u}}{2\ell}\right)\left(r\cos\dfrac{u}{\ell}+\ell\sin\dfrac{u}{\ell}\right)}{r\cos\dfrac{l_{\phi}}{2}},$
$\cos^{2}\phi=\cos^{2}\dfrac{l_{u}}{2\ell}+\dfrac{\ell^{2}\sin^{2}\dfrac{l_{u}}{2\ell}\cos^{2}\dfrac{l_{u}}{2\ell}}{r^{2}\left(\sin^{2}\dfrac{l_{u}}{2\ell}-\sin^{2}\dfrac{l_{\phi}}{2}\right)}.$ (3.10)
Note that we still have the condition $\sin\dfrac{l_{\phi}}{2}>\sin{\dfrac{l_{u}}{2\ell}}$ ) which ensures that the new curve remains spacelike. From the second equation of (3.1), we find that the radial coordinate $r$ has a minimum value which we denote as $r_{min}$ and
$r_{min}=\dfrac{\ell\sin\dfrac{l_{u}}{2\ell}}{\sqrt{\sin^{2}\dfrac{l_{\phi}}{2}-\sin^{2}\dfrac{l_{u}}{2\ell}}}.$ (3.11)
Moreover, our new extremal curve consists of two disconnected branches. Using the definition of $r_{min}$ (3.11), we can write (3.1) in its most simplified form,
$\cos\phi=\dfrac{\cos\left(\dfrac{l_{u}}{2\ell}\right)\left(r\cos\dfrac{u}{\ell}+\ell\sin\dfrac{u}{\ell}\right)}{r\cos\dfrac{l_{\phi}}{2}},$
$\cos(\phi)=\pm\dfrac{\cos\dfrac{l_{u}}{2\ell}}{r}\sqrt{r^{2}-r_{min}^{2}},$ (3.12)
where $\pm$ shows the equation of two branches. Combining two equations of (3.1) we can omit $\phi$ and find
$r\cos\frac{u}{\ell}+\ell\sin\dfrac{u}{\ell}=\pm\cos\dfrac{l_{\phi}}{2}\sqrt{r^{2}-r_{min}^{2}}.$ (3.13)
Using this equation we can determine the extremum of $u$. For both branches, these values occur at the same radial coordinate $r=r_{I}$ where
$r_{I}=\dfrac{\ell\sin\dfrac{l_{u}}{2\ell}\cos\dfrac{l_{u}}{2\ell}}{\sin^{2}\dfrac{l_{\phi}}{2}-\sin^{2}\dfrac{l_{u}}{2\ell}}.$ (3.14)
3.2 Calculation of length and taking flat-space limit
To find the lengths of the two branches, we first write them in the Poincare coordinate (3.1). From (3.1) we observe that exchanging $l_{u}\leftrightarrow\ell l_{\phi}$ in the BMS-global coordinate is equivalent to exchanging $l_{x}\leftrightarrow l_{t}$ in the Poincare coordinate. Therefore, we can write our new extremal curve (3.1) in terms of the Poincare coordinate by starting from (3.1) and performing the exchange $l_{x}\leftrightarrow l_{t}$,
$x^{2}=\dfrac{l_{t}^{2}\left(l_{t}^{2}-l_{x}^{2}-4z^{2}\right)}{4\left(l_{t}^{2}-l_{x}^{2}\right)}$
$t^{2}=\dfrac{l_{x}^{2}\left(l_{t}^{2}-l_{x}^{2}-4z^{2}\right)}{4\left(l_{t}^{2}-l_{x}^{2}\right)}.$ (3.15)
The condition $l_{x}>l_{t}$ is again imposed to ensure that this curve remains spacelike for all values of $z$. We see that where $r=r_{min}$ or equivalently $\cos\phi=0$ in the global BMS coordinates, it results in $\Delta=0$ in (3.1) which corresponds to $\{z=\infty,t=\infty,x=\infty\}$. Furthermore, $r=r_{I}$ in the global BMS coordinate maps to $z=z_{I}$, where
$z_{I}=\dfrac{2\ell^{2}}{\left(\cos\dfrac{l_{\phi}}{2}+\cos\dfrac{l_{u}}{2}\right)\sqrt{r_{I}^{2}-r_{min}^{2}}}.$ (3.16)
Using (3.1), (3.11) and (3.14), we can write (3.16) as
$z_{I}=4\ell\sqrt{64\ell^{2}l_{x}^{2}+\left(16\ell^{2}-l_{x}^{2}+l_{t}^{2}\right)^{2}}\left(\dfrac{2\ell}{l_{t}\left(16\ell^{2}-l_{x}^{2}+l_{t}^{2}\right)}-\dfrac{1}{8\ell l_{x}}\right).$ (3.17)
Using (3.1) and (3.2) we can simplify the line element as
$ds=\dfrac{\ell Adz}{z\sqrt{z^{2}+A^{2}}},$ (3.18)
where
$A=\dfrac{\sqrt{l_{x}^{2}-l_{t}^{2}}}{2}.$ (3.19)
We can use (3.1) to write (3.18) in terms of the parameters of the global BMS coordinate (3.4),
$ds=\dfrac{-\ell d\Delta}{\sqrt{\Delta^{2}+\dfrac{4\ell^{4}}{A^{2}}}}.$ (3.20)
From (3.1), (3.1) and (3.11), we see that $\Delta$ is zero when $r=r_{min}$. Thus, integrating (3.20) from an arbitrary radius $r$ corresponding to $\Delta=\Delta_{i}$ to $r=r_{min}$ or $\Delta=0$, gives the following length for the two branches:
$L=2\ell\sinh^{-1}\left(\dfrac{A\Delta_{i}}{2\ell^{2}}\right).$ (3.21)
If $\Delta_{i}$ corresponds to the boundary of spacetime, $\dfrac{L}{4G}$ is exactly the spacelike entanglement entropy of the original spacelike interval or the real part of the timelike entanglement entropy of the timelike interval. However, the flat-space limit of this quantity is not well-defined. It can be verified that a well-defined flat space limit is achievable when $\Delta_{i}$ corresponds to $r=r_{I}$ where $r_{I}$ is given by (3.14). In this case, by considering only a portion of the curves from $r_{min}$ to $r_{I}$, $\Delta_{i}=\Delta_{I}$ where
$\Delta_{I}=\dfrac{\ell\sin\dfrac{l_{u}}{2\ell}\cos\dfrac{l_{\phi}}{2}}{\cos\dfrac{l_{u}}{2\ell}-\cos\dfrac{l_{\phi}}{2}}$ (3.22)
and
$L=2\ell\sinh^{-1}\left(\dfrac{\sin\dfrac{l_{u}}{2\ell}\cos\dfrac{l_{\phi}}{2}}{\sqrt{\cos^{2}\dfrac{l_{u}}{2\ell}-\cos^{2}\dfrac{l_{\phi}}{2}}}\right)$ (3.23)
The flat-space limit of (3.23) is well-defined and corresponds exactly to the entanglement entropy of an interval in the dual theory of flat spacetime [15] . It is evident that (3.23) does not represent the entanglement entropy of an interval in the original CFT. An interpretation of this quantity in the context of the two dimensional CFT remains unresolved in this paper. Note that $L$ in (3.23) is not invariant under any boost applied to the boundary interval. For instance if $l_{u}=0$, then $L=0$, whereas for the boosted intervals $L\neq 0$
4 Holographic pseudo entropy in dS and its flat-space limit
To find extremal curves with a well-defined flat-space limit in dS spacetimes, we use the fact that the metric of dS3 can be obtained from the metric of AdS3 by replacing $\ell^{2}\to-\ell^{2}$. Thus, we start form dS3 written in the BMS coordinate as
$ds^{2}=-\left(1-\dfrac{r^{2}}{\ell^{2}}\right)du^{2}-2dudr+r^{2}d\phi^{2},$ (4.1)
where $\ell$ is the dS radius and the cosmological horizon is located at $r=\ell$. Instead of calculating extremal curves directly for this spacetime, we can use the results of previous section and replace $\ell$ with $i\ell$. However the RT curve (3.1), after this substitution, still does not yield a well-defined flat-space limit. Therefore, we need to construct an extremal curve with a well-defined flat-space limit by starting from our new curve (3.1) and substituting $\ell$ with $i\ell$. We propose the following extremal curve as the one with a well-defined flat space limit in dS spacetimes:
$r=\dfrac{\ell\cosh\left(\dfrac{l_{u}}{2\ell}\right)\sinh\left(\frac{u}{\ell}\right)}{\cos\left(\dfrac{l_{\phi}}{2}\right)\cos(\phi)-\cosh\left(\dfrac{l_{u}}{2\ell}\right)\cosh(\dfrac{u}{\ell})},$
$r^{2}\sin^{2}(\phi)\sin^{2}\left(\dfrac{l_{\phi}}{2}\right)=\sinh^{2}\left(\dfrac{l_{u}}{2\ell}\right)\left(\ell\cosh\left(u\over\ell\right)+r\sinh\left(\dfrac{u}{\ell}\right)\right)^{2}.$ (4.2)
To find the extension of this curve into the bulk, we first simplify (4) to
$\cos\phi=\dfrac{\cosh\dfrac{l_{u}}{2\ell}\left(\ell\sinh\dfrac{u}{\ell}+r\cosh\dfrac{u}{\ell}\right)}{r\cos\dfrac{l_{\phi}}{2}},$
$\cos^{2}\phi=\cosh^{2}\dfrac{l_{u}}{2\ell}-\dfrac{\ell^{2}\sinh^{2}\dfrac{l_{u}}{2\ell}\cosh^{2}\dfrac{l_{u}}{2\ell}}{r^{2}\left(\sin^{2}\dfrac{l_{\phi}}{2}+\sinh^{2}\dfrac{l_{u}}{2\ell}\right)}.$ (4.3)
From the second equation in (4) we find that $r\geq r_{min}$ where
$r_{min}=\dfrac{\ell\sinh\frac{l_{u}}{2\ell}}{\sqrt{\sinh^{2}\dfrac{l_{u}}{2\ell}+\sin^{2}\dfrac{l_{\phi}}{2}}}.$ (4.4)
It is clear that $r_{min}<\ell$. Thus this extremal curve has a turning point at $r=r_{min}$ which is inside the cosmological horizon. Using (4.4) we can write (4) as
$\cos\phi=\pm\cosh\dfrac{l_{u}}{2\ell}\sqrt{1-\dfrac{r_{min}^{2}}{r^{2}}},$
$\ell\sinh\dfrac{u}{\ell}+r\cosh\dfrac{u}{\ell}=\pm r\cos\dfrac{l_{\phi}}{2}\sqrt{1-\dfrac{r_{min}^{2}}{r^{2}}}.$ (4.5)
From the first equation in (4), we conclude that our extremal curve is truncated at a point where $r=r_{max}$ and
$r_{max}=\dfrac{\ell\cosh\dfrac{l_{u}}{2\ell}}{\sqrt{\sinh^{2}\dfrac{l_{u}}{2\ell}+\sin^{2}\dfrac{l_{\phi}}{2}}}.$ (4.6)
Note that $r_{max}>\ell$, and for our new extremal curve, $r_{min}<r<r_{max}$. This curve obviously crosses the cosmological horizon. Moreover, the flat space limit of (4) matches the results of [16]. Our new curve is precisely the spacelike part of the curve proposed in [11] as the holographic dual of pseudo entropy in dS/CFT.
From the second equation of (4), we can determine the radius at which $u$ is extremum. Denoting this radius by $r_{I}$, and setting $\dfrac{du}{dr}\Big{|}_{r=r_{I}}\!\!\!\!\!\!\!=0$, in the second equation of (4), we obtain the following interesting equation:
$r_{I}=\dfrac{r_{min}r_{max}}{\ell}.$ (4.7)
Using (4.4), (4.6) and (4.7), we find that $r_{I}\leq\ell$. Thus a portion of our extremal curve, characterized by $r_{min}\leq r\leq r_{I}$, is always inside the cosmological horizon.
To find the line element in this case, it is sufficient to replace $\ell$ with $i\ell$ in (3.20). We obtain
$ds=\dfrac{\ell d\Delta}{\sqrt{B^{2}-\Delta^{2}}},$ (4.8)
where
$\Delta=\ell\sinh\dfrac{u}{\ell}+r\cosh\dfrac{u}{\ell}+r\cos\phi,$
$B^{2}=\dfrac{\ell^{2}\left(\cos\frac{l_{\phi}}{2}+\cosh\frac{l_{u}}{2\ell}\right)^{2}}{\sin^{2}\frac{l_{\phi}}{2}+\sinh^{2}\dfrac{l_{u}}{2\ell}}.$ (4.9)
Our curve will be spacelike if $B^{2}>\Delta^{2}$. Using the equations for the curve (4) along with (4), we find that this condition is satisfied for $r<r_{max}$. Therefore, our new extremal curve in dS spacetime is always spacelike. Moreover, using the equations for the curve (4) and the definition of $r_{min}$ and $r_{max}$ in (4.4) and (4.6), we simplify (4.8) as
$ds=\dfrac{\ell rdr}{\sqrt{\left(r^{2}-r_{min}^{2}\right)\left(r_{max}^{2}-r^{2}\right)}}.$ (4.10)
Since this extremal curve consists of two similar branches, the total length, $L$ of the curve is twice the integral of (4.10) from $r_{min}$ to $r_{max}$. Evaluating this integral yields, $L=\ell\pi$. Note that $L/4G$ is half of the entropy of the cosmological horizon and is exactly the real part of the pseudo entropy calculated in [11].
While the flat-space limit of (4) is well-defined ,the length diverges as $\ell\to\infty$. This indicates that to take the flat-space limit, one must consider only a portion of this curve. We find that this portion must be characterized by $r_{min}<r<r_{I}$, where $r_{I}$ is given by (4.7). If we calculate the integral of (4.10) from $r_{min}$ to $r_{I}$, the length of the two branches will be
$L_{I}=2\ell\sin^{-1}\dfrac{\sinh\frac{l_{u}}{2\ell}\cos\frac{l_{\phi}}{2}}{\sqrt{\sinh^{2}\dfrac{l_{u}}{2\ell}+\sin^{2}\frac{l_{\phi}}{2}}}.$ (4.11)
A simple calculation gives
$\dfrac{1}{4G}\lim_{\ell\to\infty}L_{I}=\dfrac{l_{u}\cos\frac{l_{\phi}}{2}}{4G\sin\frac{l_{\phi}}{2}},$ (4.12)
which is precisely the entanglement entropy of CCFT (2.7) with $c_{L}=0$ and $c_{M}=\dfrac{3}{G}$.
5 Summary and future directions
In this paper, we demonstrate that the extremal spacelike curves in 3D flat spacetime, whose lengths correspond to the entanglement entropy in CCFT2, can be derived by taking the flat-space limit from the spacelike geodesics in dS and AdS spacetime. The length of these geodesics yields the real part of pseudo-entropy in CFT2. Extending this relation to dual field theories suggests that the entanglement entropy in CCFT can be obtained from pseudo-entropy by taking the ultra-relativistic limit from CFT. Accepting this connection raises several questions that need to be addressed in future works. One such question is how imaginary part of pseudo-entropy and its holographic dual, which are timelike geodesics, are canceled in the limiting process.
Another issue discussed in this article is that to accurately obtain the entanglement entropy in CCFT, one should calculate only the length of a segment of the spacelike curve rather than its entirety. This segment length is calculated from a point in dS or AdS spacetime where the time direction, $u$, is extremum. The reason for this and its significance in dual field theories are questions we aim to answer in future studies. This problem becomes more intriguing when we consider formulas (3.23) and (4.11), which result from calculating the segment length of spacelike curves before taking the flat-space limit. These two formulas, which can be transformed into each other by the $\ell\to i\ell$ transformation, are quantitative expressions in Lorentzian CFT and Euclidean CFT, respectively, whose flat-space limit yields the entanglement entropy of CCFT. Understanding this quantity and its relation to pseudo-entropy or entanglement entropy is also a compelling question that will be the focus of our future researches.
Acknowledgments
The author would like to thank M. Ghanbari for his contribution in the first stage of this work.

### v2

@@
-August 6, 2024
+October 6, 2024
 Abstract
 The real part of pseudo entropy in conformal field theories is holographically calculated by the area of some extremal spacelike surfaces in the dual dS and AdS spacetimes.
-We show that the flat-space limit of these curves in three dimensional (A)dS spacetimes is well-defined.
+We show that the flat-space limit of these curves in three-dimensional (A)dS spacetimes is well defined.
 We find that if the length of the curves is calculated from the radial coordinate where the retarded time is extremum, then after taking the flat-space limit, the entanglement entropy of the dual theory of three-dimensional flat spacetime is obtained.
-For dS spacetime, the radial coordinate corresponding to the extremum of retarded time is located inside the cosmlogical horizon.
-Our results suggest that on the field theory side, the entanglement entropy in the dual theory of flat spacetimes should be obtained from the ultra-relativistic limit of pseudo entropy in the dual CFT to (A)ds spacetimes.
+For dS spacetime, the radial coordinate corresponding to the extremum of retarded time is located inside the cosmological horizon.
+Our results suggest that on the field theory side, the entanglement entropy in the dual theory of flat spacetimes should be obtained from the ultrarelativistic limit of pseudo entropy in the dual CFT to (A)dS spacetimes.
 1 Introduction
-It has been proposed in [1]-[2] that the flat-space limit (i.e., the zero cosmological constant limit) of asymptotically AdS spacetimes is equivalent to the ultra-relativistic limit [3] in conformal field theory.
+It has been proposed in [1]-[2] that the flat-space limit (i.e., the zero cosmological constant limit) of asymptotically AdS spacetimes is equivalent to the ultrarelativistic limit [3] in conformal field theory.
 As a result, it has been suggested that the dual field theory to asymptotically flat spacetimes in three dimensions is a two-dimensional ultra-relativistic field theory, known as Carrollian Conformal Field Theory (CCFT) [2, 4, 5].
@@
 For CFTs, the holographic dual of this entanglement entropy, according to Ryu-Takayanagi [10], is represented by extremal spacelike surfaces in the bulk of asymptotically AdS spacetimes.
-Recently, [11]-[12] have studied the generalization of entanglement entropy for timelike subsystems and proposed holographic duals for these as well.
+Recently, the authors of [11]-[12] have studied the generalization of entanglement entropy for timelike subsystems and proposed holographic duals for these as well.
 This entropy, known as timelike entanglement entropy, is a complex quantity where the real part corresponds to the area of spacelike extremal surfaces and the imaginary part corresponds to the area of timelike extremal surfaces in the bulk.
@@
 This field theory, defined at the timelike infinity of dS spacetime, is non-unitary, so the entanglement entropy for its subsystems is essentially pseudo-entropy and must be complex.
-The holographic dual of this pseudo-entropy is given in [11]-[12], consisting of extremal surfaces in the bulk, with parts being spacelike and others timelike.
+The holographic dual of this pseudo-entropy is given in [11]-[12], consisting of extremal surfaces in the bulk, with parts being spacelike and others timelike (see also [15]-[19]).
 Unlike in AdS, the timelike surfaces here connect to the subsystem at timelike infinity, while the spacelike surfaces appear inside the bulk and connect to the timelike surfaces.
 For CCFTs, as the dual theory of asymptotically flat spacetimes, entanglement entropy can also be defined.
-Using the infinite-dimensional symmetry of two-dimensional CCFTs, a universal formula for the entanglement entropy of intervals is provided in [15].
+Using the infinite-dimensional symmetry of two-dimensional CCFTs, a universal formula for the entanglement entropy of intervals is provided in [20] (see also the recent work [21]).
 The holographic interpretation of this entropy is proposed in [22] as extremal curves in asymptotically flat spacetimes.
@@
 However, in this case, the entropy is complex-valued.
-This entropy is reffered to as timelike entanglement entropy.
+This entropy is referred to as timelike entanglement entropy.
 The complexity arises because the density matrix in this case is not Hermitian.
@@
 2.2 Entanglement entropy in the dual theory of flat spacetimes
-In paper [1]-[2] , it is proposed that the dual field theory to asymptotically flat spacetime in three dimensions is a two-dimensional CCFT, which possesses the following infinite-dimensional symmetry:
+In papers [1]-[2] , it is proposed that the dual field theory to asymptotically flat spacetime in three dimensions is a two-dimensional CCFT, which possesses the following infinite-dimensional symmetry:
 $[L_{n},L_{m}]=(n-m)L_{n+m}+\dfrac{c_{L}}{12}(n^{3}-n)\delta_{n+m,0},$
 $[L_{n},M_{m}]=(n-m)M_{n+m}+\dfrac{c_{M}}{12}(n^{3}-n)\delta_{n+m,0},$
-$[M_{n},M_{m}]=0.$ (2.5)
-$c_{L}$ and $c_{M}$ are central charges.
+$[M_{n},M_{m}]=0,$ (2.5)
+where $c_{L}$ and $c_{M}$ are central charges.
 This algebra is derived from the Inonu-Wigner contraction of the conformal algebra in two dimensions.
@@
 It is evident from algebra (2.2) that the symmetries of two-dimensional CCFTs are infinite dimensional, similar to those of two-dimensional CFTs .
-This infinite-dimensional symmetry allow for the identification of universal properties within these theories.
+This infinite-dimensional symmetry allows for the identification of universal properties within these theories.
 Among these features are universal formulas for entanglement entropy.
-Research into this aspect of two-dimensional CCFTs began in [15] , where it it was shown that, in a CCFT on the plane, if we consider the subsystem to be the interval $\left(-\dfrac{l_{u}}{2}<u<\dfrac{l_{u}}{2},\,-\dfrac{l_{\phi}}{2}<\phi<\dfrac{l_{\phi}}{2}\right)$ , the entanglement entropy is given by
+Research into this aspect of two-dimensional CCFTs began in [20], where it was shown that, in a CCFT on the plane with coordinates $(u,\phi)$ where $u$ is known as retarded time [8] and $\phi$ is non-periodic, if we consider the sub-system to be a line with end points $\left(u=-\dfrac{l_{u}}{2},\phi=-\dfrac{l_{\phi}}{2}\right)$ and $\left(u=\dfrac{l_{u}}{2},\phi=\dfrac{l_{\phi}}{2}\right)$ where $l_{u}$ and $l_{\phi}$ are constants, the entanglement entropy is given by
 $S_{EE}=\dfrac{c_{L}}{6}\log\dfrac{l_{\phi}}{\epsilon}+\dfrac{c_{M}}{6}\dfrac{l_{u}}{l_{\phi}},$ (2.6)
-where $\epsilon$ is a cutoff used to regulate the interval. For CCFTs on the cylinder this formula is given by
+where $\epsilon$ is a cutoff used to regulate the interval and $c_{L}$ and $c_{M}$ are central charges of (2.2).
+For CCFTs on the cylinder with coordinates $(u,\phi)$ where $\phi$ is periodic, this formula is given by
 $S_{EE}=\dfrac{c_{L}}{6}\log\left(\dfrac{2}{\epsilon}\sin\dfrac{l_{\phi}}{2}\right)+\dfrac{c_{M}}{12}l_{u}\cot\dfrac{l_{\phi}}{2}.$ (2.7)
@@
 This implies that the extremal spacelike curves , whose length leads to the entropies (2.6) and (2.7) , do not connect the two ends of the interval directly.
-In fact, the proposal of [16] is that the holographic interpretation in this context involves spacelike curve whose endpoints are connected to the two ends of the interval at null infinity by two null rays.
+In fact, the proposal of [22] is that the holographic interpretation in this context involves spacelike curves whose end points are connected to the two ends of the interval at null infinity by two null rays.
 3 Holographic pseudo entropy in AdS and its flat limit
@@
 The holographic dual of this entropy introduced in [11]-[12] consists of two geodesics: one spacelike and one timelike.
-To obtain the flay-space limit, we only need the spacelike part.
+To obtain the flat-space limit, we only need the spacelike part.
 We begin by considering a spacelike interval on the boundary of AdS3, compute its holographic entanglement entropy, and then perform a Wick rotation to extract the spacelike part of the holographic timelike entanglement entropy.
@@
 To find extremal curves with a well-defined flat-space limit in dS spacetimes, we use the fact that the metric of dS3 can be obtained from the metric of AdS3 by replacing $\ell^{2}\to-\ell^{2}$.
-Thus, we start form dS3 written in the BMS coordinate as
+Thus, we start from dS3 written in the BMS coordinate as
 $ds^{2}=-\left(1-\dfrac{r^{2}}{\ell^{2}}\right)du^{2}-2dudr+r^{2}d\phi^{2},$ (4.1)
@@
 These two formulas, which can be transformed into each other by the $\ell\to i\ell$ transformation, are quantitative expressions in Lorentzian CFT and Euclidean CFT, respectively, whose flat-space limit yields the entanglement entropy of CCFT.
-Understanding this quantity and its relation to pseudo-entropy or entanglement entropy is also a compelling question that will be the focus of our future researches.
-Acknowledgments
+Understanding this quantity and its relation to pseudo-entropy or entanglement entropy is also a compelling question that will be the focus of our future research.
+Acknowledgment
 The author would like to thank M. Ghanbari for his contribution in the first stage of this work.